OpenStudy (anonymous):

Solve the following equation: 4x2 + 13x = –3.

OpenStudy (anonymous):

Come on allybally, people have been showing you how to solve quadratics all day ;) \[\Large \begin{array}{l} 4{x^2} + 13x = - 3\\ 4{x^2} + 13x + 3 = 0\\ \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}} \end{array}\] Or you can also factorise it to \(\Large (4x + 1)(x + 3)\)

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

If we put x=-3 then eqn becomes 4x(-3)^2+13x(-3)=-3